What does "x" equal in these equations?
1. 2x+48=10x
2. 5x-26=18x
3. 4x+4=2x+36

Answer by user_dude2008(1862)

1.
2x+48=10x
48=10x-2x
48=8x
48/8=x
6=x
Answer: x=6

2.
5x-26=18x
-26=18x-5x
-26=13x
-26/13=x
-2=x
Answer: x=-2

3.
4x+4=2x+36
4x-2x=36-4
2x=32
x=32/2
x=16
Answer: x=16